(25 x 9) + (75 x 9) =
1
Expert's answer
2019-10-29T06:17:18-0400

Multipliers of 75 are 25 and 3. Then

"(25 * 9) + (75 * 9) =(25 * 9) + (25*3* 9)=..."

|Take a common part out the brackets|

"...=(25*9)(1+3) = 25*9*4 = ..."

|We use that "25 * 4 = 100"|

"...=100*9=900" .

So, the complete solution is

"(25 * 9) + (75 * 9) =(25 * 9) + (25*3* 9)=(25*9)(1+3) = 25*9*4 =100*9=900."

Answer: "(25 * 9) + (75 * 9) =900" .